OpenStudy (anonymous):

find the vertex, the line of symmetry, the maximum or minimum value of the quadratic function.. f(x)=3x^2-18x+1

OpenStudy (anonymous):

x value of the vertex is -b/2a = 18/6 = 3, plug that in to get y =3 *3^2-18*3+1 = -26 vertex (3,-26). The parabola opens up so -26 is the minimum. The line of symmetry is x=3